OpenStudy (anonymous):

Find the cube roots of 8(cos 216° + i sin 216°).

OpenStudy (shubhamsrg):

Euler theorem to the aid! B|

hartnn (hartnn):

8(cos 216° + i sin 216°).

OpenStudy (anonymous):

Is polar form a + bi (standard form)?

hartnn (hartnn):

can you write 8(cos 216° + i sin 216°) in polar form ? r (angle) theta

terenzreignz (terenzreignz):

Is it not already in polar form?

hartnn (hartnn):

ok, i meant 8 <216

terenzreignz (terenzreignz):

Although, for the sake of ease of typing, might we set r(cos A + isin A) = r cis A

hartnn (hartnn):

so, in such a case, to find cube root, first just take the cube root of magnitude so, sube root of 8 =... ?

hartnn (hartnn):

*cube

OpenStudy (anonymous):

is 2

hartnn (hartnn):

and just do 1/3 of the angle..

terenzreignz (terenzreignz):

Use this formula.... \[\large \left[ r(\cos \theta + i \sin \theta) \right]^p=r^p(\cos \ p \theta + i \sin \ p \theta)\]

terenzreignz (terenzreignz):

And remember that taking the cube root is just raising to a fractional exponent, 1/3.

OpenStudy (anonymous):

Oh

hartnn (hartnn):

i was typing that only.. p=1/3

OpenStudy (anonymous):

So at k=0, 2(cos pi/6 + isin pi/6)

terenzreignz (terenzreignz):

But... with fractional exponents, it's not that simple... but it isn't too complicated neither :) There's a slight twist XD

OpenStudy (anonymous):

And at k=1, 2(cos 5pi/6 + isin 5pi/6)

terenzreignz (terenzreignz):

Wait a minute, how did you get pi/6?

OpenStudy (anonymous):

@hartnn Please give @terenzreignz a medal for me! Thanks to both of you for your help, I got it. :)

hartnn (hartnn):

i gave him/her way before you thought...

terenzreignz (terenzreignz):

I think that's been done :) But where did you get pi/6 ? @hartnn I'm Terence --> I'm a boy XD

hartnn (hartnn):

noted.